OpenStudy (wasiqss):

Integrate (2z)/(5z-1)

OpenStudy (wasiqss):

plz urgent help

OpenStudy (turingtest):

um... long division?

OpenStudy (wasiqss):

its very complex by long divsionn,

OpenStudy (anonymous):

use this. http://www.wolframalpha.com/bing/?i=integral+of+(2z)%2f(5z-1)&expl=3 all answers.

OpenStudy (mr.math):

Use can put \(u=5z-1\), this is one way.

OpenStudy (wasiqss):

Mr m wrong

OpenStudy (turingtest):

\[\frac{2z}{5z-1}=\frac25+\frac25\frac1{5z-1}\]

OpenStudy (wasiqss):

turing test cant understand the form plz write it clearly

OpenStudy (turingtest):

? long division\[\int \frac{2z}{5z-1}=\int \frac25+\frac25\frac1{5z-1}dz\]\]

OpenStudy (turingtest):

I don't really know how to illustrate the process on this site...

OpenStudy (turingtest):

Mr math's sub works too

OpenStudy (mr.math):

No, it's not wrong. Have you even tried it?

OpenStudy (agreene):

you can factor the 2 in the numerator out and it will be a bit simpler: \[\int \frac{2z}{5z-1} dz= 2\int \frac{z}{5z-1}dz=2\int\frac15+\frac1{5(5z-1)}dz\] Then split the integrals and go from there.

OpenStudy (wasiqss):

thanks